The continuity equation is satisfied by *

u = A sin xy, v = - A sin xy

u = x + y. V = x - y

u = 2x2 + cy, v = 3y2

Ou = x + 2y, v = 2x + y.
1
Expert's answer
2021-07-15T10:51:59-0400

The continuity equation is satisfied by

"\text{YES}~u = A \sin xy,~ v = - A \sin xy"

"\text{YES}~u = x + y,~v = x - y"

"\text{NO}~u = 2x^2 + cy, ~v = 3y^2"

"\text{YES}~u = x + 2y, ~v = 2x + y"